OpenStudy (anonymous):

choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6).

OpenStudy (mrnood):

choose 1 point as a and the other as b - it doesn't matter which. Then calculkate ya-yb and xa-xb divide the first number you got by th second - that is th eslope. Show oyur work and I will help

OpenStudy (anonymous):

-4-8/-3--6 = -12/3=-4/1

OpenStudy (anonymous):

but -4 isnt a answer choice?

OpenStudy (mrnood):

Close - but needs another look... in the brackets the FIRST number is X and the second is Y - I think you got them mixed up.

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\)

OpenStudy (anonymous):

-6-8/-3-4=-14/-7=-2/1

OpenStudy (mrnood):